please help me. I have trouble completing the square to find the center & radius. The problem is:
x^2+y^2+2x+10y+25=0 ;
I need the answer & I need to know how to do more of those problems, thanks.

Answer by rwm(914)

Separate x's and y's and constants
This is a circle.
x^2+2x+(y+5)^2 = 0
(x+1)^2+(y+5)^2 = 1
center -1,-5
radius 1
area =pi
perimeter (circumference) =2pi